OpenStudy (anonymous):

g(x) = √5x - 4 h(x) = 4x2 + 7 Find: h(g(1))

OpenStudy (anonymous):

First find g(1) by plugging 1 in for the x in g(x)

OpenStudy (anonymous):

Then take that answer and plug it into the x in h(x)

OpenStudy (anonymous):

i tried that but it wasnt one of the choice anwers.

OpenStudy (campbell_st):

simple question is \[g(x) = \sqrt{5x - 4}\] or \[g(x) = \sqrt{5x} - 4\]

OpenStudy (anonymous):

what did you get when you plugged 1 into the g(X)